A Bibliography of Publications in Electronic Composition and Imaging

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 322 INSCC

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 January 1999 Version 1.04

Title word cross-reference

4.0 [115, 116].

6 [119].

Acrobat [118]. Adds [44]. Adobe [36, 57, 118]. Affair [4]. after [82]. Aldus [44]. Alec [92]. Alleviating [64]. ambitions [45]. Amiga [17, 24]. Annual [31, 71, 111]. Ante [23]. Applications [68]. Art [4, 59]. Arts [5]. asking [51]. Association [111]. Atari [25]. attention [118]. Awards [111]. Awareness [110].

Back [39, 85]. back-end [85]. backup [49].

Bad [102]. Basic [39]. Battling [38]. Be [1]. been [118]. below [112]. benefit [118]. Best [21]. Beta [11]. Better [26]. between [118]. Big [119]. blues [64]. Boards [9]. Boiler [112]. Book [34]. Boost [109]. BOSS [38]. Boston [103]. Brainerd [72]. Bridge [112]. brought [121]. bugs [115, 116]. Builder [4]. Bureau [12, 13, 71]. Bureaus [77, 89]. Business [56].

Can [55, 118]. Canadian [111]. capital [48]. Caps [48]. Caught [101]. CD [70, 86]. CD-ROM’s [70]. CEO [62]. Changing [12]. Chasing [74]. Choosing [18]. Clients [83]. close [118]. Color [15, 16, 18, 29, 37, 45, 96, 98, 101, 107]. colorful [45]. Come [63, 70].

Comes [10, 15, 52]. Communications [10]. company [45]. Comparisons [43]. compression [62]. Computer [57, 111]. Computers [19]. Contest [61]. Corel [58, 115, 116]. Coreldraw [16, 61, 79, 115, 116]. Corporate [69]. Correcting [84]. Couckuyt [92]. Cover [33]. Cowboy [55]. Created [20]. Creating [33, 59, 69]. creativity [112].

dancing [65]. Darkroom [30]. Data [64]. Debunking [86]. decks [112]. Densitometry [107]. Design [19, 61]. designed [118]. Designer [23, 53]. Designers [83]. Desktop [101, 104]. did [114]. Digital [30, 91, 120]. Dilemma [40]. Directory [31, 71]. discusses [62]. disk [42, 64]. Do [6, 64, 75]. Documents [120]. does [118]. Domain [8]. Don’t [117]. Double [42]. Dr. [47].

EC [31, 33, 47, 50, 72, 92, 118]. editing [52]. Editor [5]. Editorial [106]. Effects [99]. EFI [45]. Electronic [17, 97, 112]. Electrons [65]. End [85, 101]. engineering [112]. Enough [88]. Environmental [110]. Equal [20]. Error [84]. Everything [54]. eyed [86].

Family [4]. Fast [88]. FAX [67]. Feature [61]. Features [122]. Featuring [38]. File [38, 62, 73]. Files [42]. Film [90]. Finally [101]. Finding [82]. Fire [58]. Firm [31]. Font [36, 43, 60, 93, 99]. Fontographer [4]. Fonts [70]. Food [1]. Force [76]. forget [117]. format [36]. Formats [38]. formatted [118]. Foundries [63]. Four [26]. Fourth [71]. Fred [67, 74, 95, 106]. front [85]. front-end [85]. full [64]. Fundamentals [13]. Future [95].

Games [56]. Gerald [47]. Gets [87].

Getting [21, 118]. Goes [106]. Good [63]. Goodbye [22]. got [67]. Graphics [5, 111, 121]. Great [19]. Grows [24].

hard [42, 64]. hardware [105]. heads [65]. Here [89]. High [35, 98, 101, 117]. High-Quality [98]. High-Speed [117]. High-Volume [117]. House [90].

Identity [69]. If [1]. Illustrator [57].

image [52]. Imaginary [74]. Imaging [35, 118]. Import [73]. important [117]. Importer [4]. Importing [59]. Ins [46]. Interview [47, 50, 72, 92]. issues [62].

Jaggies [22].

Know [75, 91, 107]. Knowledge [81].

Kodak [86].

Laser [117]. Last [117]. latest [105]. Lazy [102]. Less [94]. Letters [5, 48]. lines [118]. Link [100]. Look [97, 118]. Looks [95]. lot [112, 118]. Love [1]. Lucky [68]. Luke [76].

major [115, 116]. Management [42, 93]. Master [36]. materials [59]. Me [66]. Media [74]. Medium [70]. Messages [84]. Midi [6]. Missing [100]. Modem [1]. Moments [19]. Mr. [7]. Multiple [36]. Multiplying [36]. Murch [47]. Music [1, 7, 25]. My [96]. mysteries [38].

Need [91, 107]. Needs [77]. networks [118]. Never [48, 88]. Next [57, 118]. Nightmare [60]. Nine [68]. Number [68].

Oblivion [65]. observations [49].

online [51]. Opposing [50]. option [117]. Other [59]. Output [73, 117]. Oxymoron [48]. pages [118]. Paper [81]. Part [71]. Paul [53, 72]. PC [2, 8, 15, 70, 98]. Perfect [94]. Performance [2, 35, 109]. Photo [86]. Photography [91]. Planning [32]. platforms [118]. Please [7]. Pleasure [42]. Plenty [51]. Plug [46]. Plug-Ins [46]. Portability [45]. Possibilities [36]. PostScript [27, 40, 68, 79]. Powerful [52]. practical [49]. Prepress [17]. President [62]. Printer [18, 68, 80]. Printers [117]. printing [75]. Prize [111]. Problems [73]. Production [34]. Productivity [85]. program [115, 116]. Public [8]. Publishing [8]. Pushing [2].

Quality [98]. quirks [115, 116].

Rainbows [74]. Re [6, 59]. Re-creating [59]. read [65]. read-write [65]. Real [87]. reference [59]. Resolution [3]. Retrofitting [27]. Returns [58]. Reviews [102]. Revolution [3]. Right [18]. Role [12]. ROM’s [70]. Room [112].

sail [82]. Says [74]. Scan [21, 26, 66]. Scanners [20, 50]. Scanning [32, 107]. season [121]. Sending [113]. Serious [56]. Service [12, 13, 71, 77, 89]. Seybold [103]. Shading [37]. share [118]. shoals [82]. should [48]. show [121]. Site [11]. Small [45, 63]. smooth [82]. Software [46, 99, 105]. Solved [73]. Some [49, 104]. Sources [59]. Spanning [14]. Special [61, 99]. Spectrum [14]. Speed [117]. Spinning [65]. Stakes [45]. Starry [86]. Starry-eyed [86]. Stay [89]. Step [117, 118]. Steps [26]. still [112]. Stop [74]. Storm [62]. strategies [42]. Strategy [108]. Street [60]. strengths [115, 116]. Studio [19]. Success [108]. Survival [90]. Sych [53].

Systems [36].

takes [118]. Tales [34]. Technology [62]. telephone [118]. There [112]. Things [63]. Time [70]. Timing [54]. Tips [119]. Top [111]. Top-Prize [111]. Toys [105]. tragedies [104]. Trainers [28]. Training [28, 31]. Tricks [119]. triumphs [104]. Two [42, 50]. Type [51, 53, 94, 119]. Typographic [36, 48].

ultimate [121]. upgrade [115, 116]. Ups [23]. Use [48, 76]. Users [102, 118]. Using [16].

vector [59]. Views [50]. Virtual [80]. Vision [86]. Volume [32, 117]. VR [87].

War [122]. weakness [115, 116].

Welcome [104]. WFN [38]. Who [77]. Window [96]. Windows [10, 52, 89, 93]. Winners [61, 111]. Work [53, 98, 118]. World [61]. write [65].

Xpress [78]. Xtending [41].

yours [51]. Yourself [41, 78].

Zen [29]. Zippity [64].

References

Ander:1990:MIM

[1] Chuck Ander. Music if music be the food of love–modem on. Electronic Composition and Imaging, 4(??):18–??, May 1990. ISSN 0838-9535.

Shilling:1990:PPP

[2] Gary Shilling. Pushing PC performance. Electronic Composition and Imaging, 4 (??):26–??, May 1990. ISSN 0838-9535.

Reis:1990:RR

[3] Charlie Reis. The resolution revolution. Electronic Composition and Imaging, 4 (??):34–??, May 1990. ISSN 0838-9535.

Dako:1990:FAF

[4] Peter Dako. A family affair: Fontographer, The Art Importer, Family Builder. Electronic Composition and Imaging, 4 (??):44–??, May 1990. ISSN 0838-9535.

Shilling:1990:ALG

[5] Gary Shilling. Arts and Letters Graphics Editor. Electronic Composition and Imaging, 4(??):48–??, May 1990. ISSN 0838-9535.

Fleming:1990:DRM

[6] Sherwood Fleming. Do re Midi. Electronic Composition and Imaging, 4(1): 12–??, March 1990. ISSN 0838-9535.

Dako:1990:MMP

[7] Peter Dako. Mr. Music please. Electronic Composition and Imaging, 4(1): 15–??, March 1990. ISSN 0838-9535.

Ander:1990:PDP

[8] Chuck Ander. Public domain PC publishing. Electronic Composition and Imaging, 4(1):38–??, March 1990. ISSN 0838-9535.

Ander:1990:MB

[9] Chuck Ander. More boards. Electronic Composition and Imaging, 4(1):41–??, March 1990. ISSN 0838-9535.

Dickman:1990:CCW

[10] Chris Dickman. Communications comes to Windows. Electronic Composition and Imaging, 4(1):43–??, March 1990. ISSN 0838-9535.

Trusty:1990:BS

[11] Patricia Trusty. Beta site. Electronic Composition and Imaging, 4(1):49–??, March 1990. ISSN 0838-9535.

McIlroy:1990:SBC

[12] Thad McIlroy. The service bureau’s changing role. Electronic Composition and Imaging, 4(1):64–??, March 1990. ISSN 0838-9535.

Mothersill:1990:SBF

[13] Gillian Mothersill. Service bureau fundamentals. Electronic Composition and Imaging, 4(1):78–??, March 1990. ISSN 0838-9535.

Wasson:1990:SS

[14] Gregory Wasson. Spanning the spectrum. Electronic Composition and Imaging, 4(3):26–??, June 1990. ISSN 0838-9535.

Anonymous:1990:CCP

[15] Anonymous. Color comes to the PC. Electronic Composition and Imaging, 4 (3):32–??, June 1990. ISSN 0838-9535.

Dickman:1990:UCC

[16] Chris Dickman. Using color in Corel-Draw. Electronic Composition and Imaging, 4(3):36–??, June 1990. ISSN 0838-9535.

Evans:1990:EPA

[17] Jeff Evans. Electronic prepress on the Amiga. Electronic Composition and Imaging, 4(3):42–??, June 1990. ISSN 0838-9535.

Reis:1990:CRC

[18] Charles Reis. Choosing the right color printer. Electronic Composition and Imaging, 4(3):48–??, June 1990. ISSN 0838-9535.

Graff:1990:GMC

[19] Tom Graff. Great moments of computers in the design studio. Electronic Composition and Imaging, 4(3):54–??, June 1990. ISSN 0838-9535.

Reis:1990:ASC

[20] Charlie Reis. Not all scanners are created equal. Electronic Composition and Imaging, 4(4):36–??, September 1990. ISSN 0838-9535.

Roth:1990:GBS

[21] Steve Roth. Getting the best scan. Electronic Composition and Imaging, 4(4): 40–??, September 1990. ISSN 0838-9535.

Shilling:1990:GJ

[22] Gary Shilling. Goodbye jaggies. Electronic Composition and Imaging, 4(4): 44–??, September 1990. ISSN 0838-9535.

Kieran:1990:DUA

[23] Michael Kieran. Designer 3 ups the ante. Electronic Composition and Imaging, 4 (4):49–??, September 1990. ISSN 0838-9535.

Dickman:1990:AG

[24] Chris Dickman. The Amiga grows up. Electronic Composition and Imaging, 4 (4):55–??, September 1990. ISSN 0838-9535.

Sauve:1990:MA

[25] Steve Sauve. Music on the Atari. Electronic Composition and Imaging, 4(4): 58–??, September 1990. ISSN 0838-9535.

Bingham:1991:FSB

[26] G. Bingham and A. Stephen. Four steps to a better scan. Electronic Composition and Imaging, 5(3):20–??, May 1991. ISSN 0838-9535.

Dickman:1991:RP

[27] Chris Dickman and Dezso Nagy. Retrofitting PostScript. Electronic Composition and Imaging, 5(3):26–??, May 1991. ISSN 0838-9535.

Bosworth:1991:TT

[28] David Bosworth. Trainers on training. Electronic Composition and Imaging, 5 (3):32–??, May 1991. ISSN 0838-9535.

Rowley:1991:ZC

[29] Margo Rowley. The zen of color. Electronic Composition and Imaging, 5(3): 38–??, May 1991. ISSN 0838-9535.

Dako:1991:DD

[30] Peter Dako. Digital darkroom 2.0. Electronic Composition and Imaging, 5(3): 42–??, May 1991. ISSN 0838-9535.

Anonymous:1991:EAT

[31] Anonymous. The EC and I Annual Training Firm Directory. Electronic Composition and Imaging, 5(3):54–??, May 1991. ISSN 0838-9535.

Rowley:1991:PVS

[32] Margo Rowley. Planning for volume scanning. Electronic Composition and Imaging, 5(5):26–??, September 1991. ISSN 0838-9535.

Shakespeare:1991:CEC

[33] Courtland Shakespeare. Creating an EC and I cover. Electronic Composition and Imaging, 5(5):30–??, September 1991. ISSN 0838-9535.

Dickman:1991:TBP

[34] Chris Dickman and Sherwood Fleming. Tales of book production. Electronic Composition and Imaging, 5(5):32–??, September 1991. ISSN 0838-9535.

Kieran:1992:HPI

[35] Michael Kieran. High performance imaging. Electronic Composition and Imaging, 6(??):6–??, June 1992. ISSN 0838-9535.

Dean:1992:MYT

[36] David Dean. Multiplying your typographic possibilities Adobe Systems’ Multiple Master font format. Electronic Composition and Imaging, 6(??):8–??, March 1992. ISSN 0838-9535.

Shakespeare:1992:CS

[37] Courtland Shakespeare. Color shading. Electronic Composition and Imaging, 6 (??):8–??, October 1992. ISSN 0838-9535.

Schettini:1992:BFF

[38] Stephen Schettini. Battling file formats featuring the mysteries of WFN BOSS. Electronic Composition and Imaging, 6 (??):12–??, March 1992. ISSN 0838-9535.

Weibel:1992:BB

[39] Bob Weibel. Back to Basic. Electronic Composition and Imaging, 6(??):12–??, June 1992. ISSN 0838-9535.

Shubitz:1992:PD

[40] Steve Shubitz. The PostScript dilemma. Electronic Composition and Imaging, 6 (??):14–??, October 1992. ISSN 0838-9535.

Buckingham:1992:XY

[41] William Buckingham. Xtending yourself. Electronic Composition and Imaging, 6(??):16–??, June 1992. ISSN 0838-9535.

Shubitz:1992:DYP

[42] Steve Shubitz. Double your pleasure, double your files two hard disk management strategies. Electronic Composition and Imaging, 6(??):18–??, March 1992. ISSN 0838-9535.

Dean:1992:FC

[43] David Dean. Font comparisons. Electronic Composition and Imaging, 6(??): 20–??, October 1992. ISSN 0838-9535.

Dako:1992:AA

[44] Peter Dako. Aldus adds up. Electronic Composition and Imaging, 6(??):22–??, June 1992. ISSN 0838-9535.

Kieran:1992:ESC

[45] Michael Kieran. EFI stakes out color portability A small company with colorful ambitions. Electronic Composition and Imaging, 6(??):22–??, March 1992. ISSN 0838-9535.

Kieran:1992:SP

[46] Michael Kieran. Software plug-ins. Electronic Composition and Imaging, 6(??): 26–??, June 1992. ISSN 0838-9535.

Anonymous:1992:EID

[47] Anonymous. The EC and I interview: Dr. Gerald Murch. Electronic Composition and Imaging, 6(??):28–??, August 1992. ISSN 0838-9535.

Haley:1992:ACT

[48] Allen Haley. All caps: A typographic oxymoron: Why you should never use all capital letters. Electronic Composition and Imaging, 6(??):28–??, October 1992. ISSN 0838-9535.

Lovaas:1992:SPO

[49] Maggie Lovaas and Felix Kramer. Some practical observations on backup. Electronic Composition and Imaging, 6(??): 30–??, March 1992. ISSN 0838-9535.

Olson:1992:EIT

[50] SCOTT Olson. The EC and I interview: Two opposing views of scanners. Electronic Composition and Imaging, 6(??): 32–??, October 1992. ISSN 0838-9535.

Dako:1992:PTY

[51] Peter Dako. Plenty of type, yours for the online asking. Electronic Composition and Imaging, 6(??):34–??, March 1992. ISSN 0838-9535.

Kieran:1992:PIE

[52] Michael Kieran. Powerful image editing comes to Windows. Electronic Composition and Imaging, 6(??):36–??, March 1992. ISSN 0838-9535.

Olson:1992:PST

[53] Scott Olson. Paul Sych: a type designer at work. Electronic Composition and Imaging, 6(??):36–??, October 1992. ISSN 0838-9535.

Olson:1992:TIE

[54] Scott Olson. Timing is everything. Electronic Composition and Imaging, 6(??): 38–??, August 1992. ISSN 0838-9535.

Dako:1992:CC

[55] PETER Dako. Cowboy in a can. Electronic Composition and Imaging, 6(??): 42–??, October 1992. ISSN 0838-9535.

McGuire:1992:GSB

[56] Mark McGuire. Games are serious business. Electronic Composition and Imaging, 6(??):42–??, August 1992. ISSN 0838-9535.

Sherman:1992:AIN

[57] Lee Sherman. Adobe Illustrator on the NeXT computer. Electronic Composition and Imaging, 6(??):42–??, March 1992. ISSN 0838-9535.

Rimmer:1992:CRF

[58] Steve Rimmer. Corel returns fire. Electronic Composition and Imaging, 6(??): 46–??, August 1992. ISSN 0838-9535.

Shakespeare:1992:ROS

[59] Courtland Shakespeare. Re-creating from other sources importing vector art reference materials. Electronic Composition and Imaging, 6(??):49–??, March 1992. ISSN 0838-9535.

Schettini:1992:NFS

[60] Stephen Schettini. Nightmare on Font Street. Electronic Composition and Imaging, 6(??):50–??, October 1992. ISSN 0838-9535.

Anonymous:1992:SFC

[61] Anonymous. Special feature: Coreldraw World Design Contest winners. Electronic Composition and Imaging, 6(??): 52–??, August 1992. ISSN 0838-9535.

Dickman:1992:PCS

[62] Chris Dickman. The President and CEO of Storm Technology discusses file compression issues. Electronic Composition and Imaging, 6(??):52–??, March 1992. ISSN 0838-9535.

Will-Harris:1992:GTC

[63] Daniel Will-Harris. Good things come from small foundries. Electronic Composition and Imaging, 6(??):54–??, October 1992. ISSN 0838-9535.

Will-Harris:1992:ZDD

[64] Daniel Will-Harris. Zippity do data alleviating the full hard disk blues. Electronic Composition and Imaging, 6(??): 54–??, March 1992. ISSN 0838-9535.

Chato:1992:SOE

[65] Paul Chato. Spinning into oblivion electrons dancing on your read-write heads. Electronic Composition and Imaging, 6 (??):57–??, March 1992. ISSN 0838-9535.

Dako:1992:SM

[66] Peter Dako. Scan by me. Electronic Composition and Imaging, 6(??):58–??, August 1992. ISSN 0838-9535.

Roth:1992:FGF

[67] Steve Roth. Fred’s got the FAX. Electronic Composition and Imaging, 6(??): 78–??, October 1992. ISSN 0838-9535.

Shubitz:1993:NIL

[68] Steve Shubitz. Nine is a lucky number: Applications for the PostScript printer. Electronic Composition and Imaging, 7 (1):8–??, January 1993. ISSN 0838-9535.

Will-Harris:1993:CCI

[69] Daniel Will-Harris. Creating a corporate identity. Electronic Composition and Imaging, 7(1):12–??, January 1993. ISSN 0838-9535.

Dean:1993:NMP

[70] David Dean. A new medium for PC fonts: CD-ROM’s time has come. Electronic Composition and Imaging, 7(1): 14–??, January 1993. ISSN 0838-9535.

ECI:1993:FAE

[71] EC and I Staff. The fourth annual EC and I service bureau directory, part I. Electronic Composition and Imaging, 7 (1):20–??, January 1993. ISSN 0838-9535.

Shilling:1993:EIP

[72] Gary Shilling. The EC and I interview: Paul Brainerd. Electronic Composition and Imaging, 7(1):38–??, January 1993. ISSN 0838-9535.

Schettini:1993:FIO

[73] Stephen Schettini. File import and output problems solved. Electronic Composition and Imaging, 7(1):42–??, January 1993. ISSN 0838-9535.

Roth:1993:IMF

[74] Steve Roth. Imaginary media: Fred says to stop chasing rainbows. Electronic Composition and Imaging, 7(1): 54–??, January 1993. ISSN 0838-9535.

Anonymous:1993:DYK

[75] Anonymous. Do you know printing? Electronic Composition and Imaging, 7 (2):6–??, March 1993. ISSN 0838-9535.

Shakespeare:1993:UFL

[76] Courtland Shakespeare. Use the force, Luke. Electronic Composition and Imaging, 7(2):8–??, March 1993. ISSN 0838-

Dean:1993:WNS

[77] David Dean. Who needs service bureaus? Electronic Composition and Imaging, 7(2):12–??, March 1993. ISSN 0838-9535.

Shubitz:1993:XY

[78] Steve Shubitz. Xpress yourself! Electronic Composition and Imaging, 7(2): 18–??, March 1993. ISSN 0838-9535.

Rimmer:1993:CP

[79] Steve Rimmer. CorelDraw and PostScript. Electronic Composition and Imaging, 7(2):22–??, March 1993. ISSN 0838-9535.

Olson:1993:VP

[80] Scott Olson. The virtual printer. Electronic Composition and Imaging, 7(2): 26–??, March 1993. ISSN 0838-9535.

Anonymous:1993:PK

[81] Anonymous. Paper knowledge. Electronic Composition and Imaging, 7(2): 28–??, March 1993. ISSN 0838-9535.

Anonymous:1993:FNS

[82] Anonymous. Finding new shoals after a smooth sail. Electronic Composition and Imaging, 7(2):30–??, March 1993. ISSN 0838-9535.

Will-Harris:1993:DC

[83] Daniel Will-Harris. Designers and clients. Electronic Composition and Imaging, 7(2):32–??, March 1993. ISSN 0838-9535.

Schettini:1993:CEM

[84] Stephen Schettini. Correcting the error messages. Electronic Composition and Imaging, 7(2):36–??, March 1993. ISSN 0838-9535.

Anonymous:1993:PFB

[85] Anonymous. Productivity from frontend to back-end. Electronic Composition and Imaging, 7(2):42–??, March 1993. ISSN 0838-9535.

Roth:1993:DKS

[86] Steve Roth. Debunking Kodak’s starryeyed vision for Photo CD. Electronic Composition and Imaging, 7(2):54–??, March 1993. ISSN 0838-9535.

Shakespeare:1993:VGR

[87] Courtland Shakespeare. VR gets real. Electronic Composition and Imaging, 7 (3):8–??, May 1993. ISSN 0838-9535.

Shubitz:1993:FIN

[88] Steve Shubitz. Fast is never fast enough. Electronic Composition and Imaging, 7 (3):12–??, May 1993. ISSN 0838-9535.

Schettini:1993:SBW

[89] Stephen Schettini. Service bureaus: Windows is here to stay. Electronic Composition and Imaging, 7(3):16–??, May 1993. ISSN 0838-9535.

Mills:1993:SFH

[90] JULIAN Mills. Survival of the film house. Electronic Composition and Imaging, 7(3):20–??, May 1993. ISSN 0838-9535.

Stewart:1993:DPW

[91] Maureen Stewart. Digital photography: What you need to know. Electronic Composition and Imaging, 7(3):24–??, May 1993. ISSN 0838-9535.

Olson:1993:EIA

[92] Scott Olson. The EC and I interview: Alec Couckuyt. Electronic Composition and Imaging, 7(3):30–??, May 1993. ISSN 0838-9535.

Dean:1993:WFM

[93] David Dean. Windows font management. Electronic Composition and Imaging, 7(3):32–??, May 1993. ISSN 0838-9535.

Will-Harris:1993:LPT

[94] Daniel Will-Harris. Less perfect type. Electronic Composition and Imaging, 7 (3):36–??, May 1993. ISSN 0838-9535.

Roth:1993:FLF

[95] Steve Roth. Fred looks to the future. Electronic Composition and Imaging, 7 (3):46–??, May 1993. ISSN 0838-9535.

Anonymous:1993:MWC

[96] Anonymous. Out my window—why color? Electronic Composition and Imaging, 7(4):6–??, July 1993. ISSN 0838-9535.

Shakespeare:1993:EL

[97] Courtland Shakespeare. The electronic look. Electronic Composition and Imaging, 7(4):8–??, July 1993. ISSN 0838-9535.

Schettini:1993:HWH

[98] Stephen Schettini. How to work with high-quality color on the PC. Electronic Composition and Imaging, 7(4):14–??, July 1993. ISSN 0838-9535.

Dean:1993:SEF

[99] David Dean. Special effects font software. Electronic Composition and Imaging, 7(4):16–??, July 1993. ISSN 0838-9535.

Shubitz:1993:ML

[100] Steve Shubitz. The missing link. Electronic Composition and Imaging, 7(4): 18–??, July 1993. ISSN 0838-9535.

Mcilroy:1993:DCF

[101] Thad Mcilroy. Has desktop color finally caught the high end? Electronic Composition and Imaging, 7(4):22–??, July 1993. ISSN 0838-9535.

Will-Harris:1993:BRL

[102] Daniel Will-Harris. Bad reviews, lazy users. Electronic Composition and Imaging, 7(4):32–??, July 1993. ISSN 0838-9535.

Schettini:1993:SB

[103] Stephen Schettini. Seybold Boston 1993. Electronic Composition and Imaging, 7 (4):33–??, July 1993. ISSN 0838-9535.

Anonymous:1993:WDT

[104] Anonymous. Welcome to desktop — some triumphs and some tragedies. Electronic Composition and Imaging, 7(4): 36–??, July 1993. ISSN 0838-9535.

Anonymous:1993:NTL

[105] Anonymous. New toys — the latest hardware and software. Electronic Composition and Imaging, 7(4):38–??, July 1993. ISSN 0838-9535.

Roth:1993:FGE

[106] Steve Roth. Fred goes editorial. Electronic Composition and Imaging, 7(4): 46–??, July 1993. ISSN 0838-9535.

Stewart:1993:CSW

[107] John Stewart. Color scanning: Why you need to know densitometry. Electronic Composition and Imaging, 7(5): 24–??, September 1993. ISSN 0838-9535.

Olson:1993:SS

[108] Scott Olson. A strategy for success.

Electronic Composition and Imaging, 7 (5):26–??, September 1993. ISSN 0838-9535.

Schettini:1993:BYP

[109] Stephen Schettini. Boost your performance. Electronic Composition and Imaging, 7(5):30–??, September 1993. ISSN 0838-9535.

Anonymous:1993:EA

[110] Anonymous. Environmental awareness.

Electronic Composition and Imaging, 7 (6):6–??, November 1993. ISSN 0838-9535.

Anonymous:1993:TWC

[111] Anonymous. The top-prize winners:

The Canadian Computer Graphics Association’s annual computer graphics awards. Electronic Composition and Imaging, 7(6):10–??, November 1993. ISSN 0838-9535.

Shakespeare:1993:BBR

[112] Courtland Shakespeare. Bridge to boiler room: There’s still a lot of engineering below the decks of electronic creativity. Electronic Composition and Imaging, 7 (6):12–??, November 1993. ISSN 0838-9535.

Dean:1993:SI

[113] David Dean. Sending it out. Electronic Composition and Imaging, 7(6):16–??, November 1993. ISSN 0838-9535.

Anonymous:1993:HWD

[114] Anonymous. How we did it. Electronic Composition and Imaging, 7(6):18–??, November 1993. ISSN 0838-9535.

Schettini:1993:CCA

[115] Stephen Schettini. Corel Coreldraw! 4.0: All the bugs, quirks, strengths and weakness of this major program upgrade. Electronic Composition and Imaging, 7 (6):20–??, November 1993. ISSN 0838-9535.

Schettni:1993:CCA

[116] Stephen Schettni. Corel Coreldraw! 4.0: All the bugs, quirks, strengths and weakness of this major program upgrade. Electronic Composition and Imaging, 7 (6):20–??, November 1993. ISSN 0838-9535.

Olson:1993:HHL

[117] Scott Olson. High-speed, high-volume laser printers: The last step. don’t forget this important output option. Electronic Composition and Imaging, 7(6): 24–??, November 1993. ISSN 0838-9535.

Rosen:1993:AAN

[118] David Rosen. Adobe Acrobat: The next step in imaging: Acrobat, designed to share formatted pages over networks, telephone lines and between platforms, has been getting a lot of attention. But how does it work? How can users benefit? EC and I takes a close look. Electronic Composition and Imaging, 7(6): 28–??, November 1993. ISSN 0838-9535.

Shubitz:1993:TTT

[119] Steve Shubitz. Type tips and tricks with the Big 6. Electronic Composition and Imaging, 7(6):30–??, November 1993. ISSN 0838-9535.

Will-Harris:1993:DD

[120] Daniel Will-Harris. Digital documents.

Electronic Composition and Imaging, 7 (6):34–??, November 1993. ISSN 0838-9535.

Anonymous:1993:WUG

[121] Anonymous. What the ultimate graphics show season brought us. Electronic Composition and Imaging, 7(6):44–??, November 1993. ISSN 0838-9535.

Roth:1993:FW

[122] Steve Roth. Features war. Electronic Composition and Imaging, 7(6):46–??, November 1993. ISSN 0838-9535.